OpenStudy (firejay5):

A 500 kg car is pushed from rest with a force of 400 N, if the car reaches a final velocity of 10 m/s, how far is it pushed? A. Find the intitial Kinetic Energy. B. Find the work done on the car. C. Find the final Kinetic Energy. D. Find the distance the car moves. (A - D ARE NOT multiple choice)

OpenStudy (shamim):

Do u know the formula F=ma

OpenStudy (firejay5):

yess @shamim

OpenStudy (shamim):

By using this formula Find a=?

OpenStudy (shamim):

If u need any more explanation plz ask me

OpenStudy (shamim):

Ok. Given mass m=500kg Right?

OpenStudy (shamim):

Force F=400N Right?

OpenStudy (danjs):

givens: mass = m = 500kg initial velocity = Vo = 0 m/s final velocity = Vf = 10 m/s Force = F = 400 N

OpenStudy (danjs):

Find the acceleration using Force = Mass x Acceleration, Acceleration = F/m then use the the attached picture to solve for the distance traveled \[\Delta x = x - x _{0}\]

OpenStudy (danjs):

Work = Force*distance

OpenStudy (danjs):

Kinetic energy = 1/2 x Mass x Velocity^2